the smaller of two consecutive odd numbers is n ; the sum of the larger and twice the smaller is 23, find the consecutive numbers
Answer by mananth(16949)

n, n+2
n+2 + 2n = 23
3n+2=23
3n=23-2
3n=21
/3
n=7
7,9